OpenStudy (anonymous):

Last year, Frank earned $7.48 per hour. This year the company decided to give him a 30% increase in his hourly rate. How much is Frank earning per hour this year? A $7.78 B $9.72 C $10.48 D $22.44 E $24.93 @MissSmartiez

OpenStudy (misssmartiez):

Construct to decimal. 0.30 x 7.48 However, I need to know. Are you finding out the balance, simple interest rate, or compound interest?

OpenStudy (anonymous):

balance

OpenStudy (misssmartiez):

ok B = P ( 1 + R) ^T

OpenStudy (misssmartiez):

7.48 * (1 + 0.30) ^ 1 year. PEMDAS 1.30^1 = 1.30. 1.30 * 7.48 = ?

OpenStudy (misssmartiez):

9.724. Round it ;) 9.72. Your answer is B, see what we did there? Tell me if you are still confused.

OpenStudy (anonymous):

9.72

OpenStudy (anonymous):

I'm still a little confused

OpenStudy (misssmartiez):

The formula is Principal, ( 1 + interest rate)^ time/year.

OpenStudy (misssmartiez):

Principal is usually the money he earns now. Rate is usually the percent. Time is hours, years, months, (mostly years)

OpenStudy (misssmartiez):

@flowercrownmgc